Calculus II Spring 2002 MA114:004–006 Russell Brown

• Homework #12 §10.1 #1, 3, 9, 11, 13, 15, 17, 19, 21, 29, 37, 51, 53, 55, 61.

• The main new idea in section 10.1 is the convergence of monotonic sequences. The study of limits of sequences should not be completely unfamiliar to students who have taken limits of functions.

• Homework #12. §10.2 #3, 9, 11, 21, 27, 29, 43, 45, 47, 55, 59. §10.3 #1, 3, 5, 7, 9, 13, 15, 19, 23, 25, 27, 29.

• In §10.2, we study two series that can be summed explicitly: geometric series and collapsing series. The geometric series is the most important series we will study.

• In section 10.3, we begin the study of tests for convergence. You should notice that every good test for convergence also gives a way to estimate the difference between the sum of a series and the partial sums. Thus, if you are interested in approximating a (hard-to-compute) series with a (easy-to-compute) partial sum, then the tests for convergence will help us to estimate the error.

• Homework G. Due Wednesday, 20 March 2002. Section 10.2, #54.

• Quiz 7. Thursday, 21 March 2002. Sections 10.2.

• There will be two more Math and Pizza! extravaganzas this semester. On Wednesday, 6 March 2002, at 4pm in CB331 Jody Fast will talk about the actuarial profession. On Wednesday, 27 March 2002, at 4pm in Chemistry-Physics 155, Frank Morgan of Williams College will talk about the mathematics of soap bubbles.

March 4, 2002